# Answers to frequently asked questions about the pulsar timing array Hellings and Downs curve

### v2

@@ Abstract
We answer frequently asked questions (FAQs) about the Hellings and Downs correlation curve—the “smoking-gun” signature that pulsar timing arrays (PTAs) have detected gravitational waves (GWs). Many of these questions arise from inadvertently applying intuition about the effects of GWs on LIGO-like detectors to the case of pulsar timing, where not all of that intuition applies. This is because Earth-based detectors, like LIGO and Virgo, have arms that are short (km scale) compared to the wavelengths of the GWs that they detect ($\approx 10^{2}$$10^{4}$ km). In contrast, PTAs respond to GWs whose wavelengths (tens of light-years) are much shorter than their arms (a typical PTA pulsar is hundreds to thousands of light-years from Earth). To demonstrate this, we calculate the time delay induced by a passing GW along an Earth-pulsar baseline (a “one-arm, one-way” detector) and compare it in the “short-arm” (LIGO-like) and “long-arm” (PTA) limits. This provides qualitative and quantitative answers to many questions about the Hellings and Downs curve. The resulting FAQ sheet should help in understanding the “evidence for GWs” recently announced by several PTA collaborations.
@@ 1 Introduction
These observations can be challenging to understand, in part because the observed signal in the PTA band is unlike the short duration ($\sim\!0.25~{}{\rm s}$) GW “chirp” signal that LIGO observed on 14 Sep 2015, from the final inspiral and merger of two $\sim\!30M_{\odot}$ black holes. Instead of short (1 s) audio frequency (100 Hz) deterministic GW signals, PTAs search for persistent (signals with correlation times $\sim\!10^{3}$ yr which are in band for $\sim\!10^{6}$ yr), low-frequency (nHz) stochastic GW signals. These are too feeble to observe directly, but should produce correlated perturbations in the arrival times of pulses from an array of galactic millisecond pulsars.
Although the current PTA observations are not statistically significant enough to unambiguously claim a detection or to identify a source, a credible candidate is the superposition of GWs from hundreds of thousands of supermassive ($\sim\!10^{9}M_{\odot}$) black holes (SMBH) orbiting one another in the centers of merging galaxies. While any one of those binaries would produce a deterministic almost-constant-frequency signal, the incoherent sum of the GWs from many such sources creates a stochastic confusion noise: a “hubbub” of GWs.
PTAs search for correlations that are predicted to follow the so-called “Hellings and Downs” curve. This correlation curve is named after Ron Hellings and George Downs, who first derived it in 1983 [5]. The idea is that pulsars act as high-quality clocks, producing radio pulses that arrive at Earth [^1] Our treatment is simplified: it assumes that the Earth is isolated, and it is not orbiting the Sun, orbited by the Moon, or affected by other planets. In practice, all pulse arrival times are converted to pulse arrival times at the center of gravity of the solar system, which is called the Solar System Barycenter (SSB). The SSB is located near the surface of the Sun, about 8 light-minutes from Earth. On the scale of the GWs of interest, this is close enough to the Earth that we do not make a distinction here. with very predictable and regular arrival times. GWs induce slight perturbations in these arrival times, and these perturbations (called timing residuals) are correlated between different pulsars in a very specific fashion.
@@ 4. Q4:
    Why isn’t the most negative value of the Hellings and Downs curve at $90^{\circ}$?
@@ 8. Q8:
    Assuming noise-free observations, are the fluctuations away from the expected Hellings and Downs curve due solely to the pulsar-term contributions to the timing residual response?
@@ 9. Q9:
Note that, as suggested by Q7, we make a distinction between (a) the Hellings and Downs curve, which is the black curve shown in Figure 1 and (b) the observed Hellings and Downs correlation, which might or might not follow this curve.
@@ 1.2 Outline
Readers who are already familiar with pulsar timing and the definition of the Hellings and Downs correlation could skip Sections 2 and 3 and jump directly to Section 4. However, we advise against that for two reasons: (i) In Sections 2.2 and 2.4, we compare the response of “long-arm” PTA pulsars to that of “short-arm” LIGO-like detectors, which we believe is the source of many of the misconceptions regarding the Hellings and Downs curve. (ii) Our discussion of pulsar averaging in Section 3.2 will be new to many, and the concepts introduced there are important for understanding what follows.
@@
From here forward, we use units in which the propagation speed $c$ of light and of GWs is equal to one. This means that both time and distance are measured in units of seconds or years as appropriate. More precisely, a distance is given by the time that it takes light to traverse that distance. For example, a meter stick is approximately 3 nanosec long, the LIGO detectors’ arms are about 10 microseconds long, the wavelengths of the GWs that PTAs are sensitive to are typically 10 years long, and typical pulsars are thousands of years from Earth.
@@ 2.1 Single GW point source
Figure 2: Geometry for the single pulsar analysis. The Earth is located at the origin of a rectangular coordinate system. The components $h_{ij}(t,\vec{x})$ of the metric perturbations are functions of time and of the Cartesian spatial coordinates $\vec{x}$. A single point source of GWs is located a distance $d$ from Earth in direction $-\hat{\Omega}$ on the sky. A galactic pulsar is located a distance $L$ from Earth in direction $\hat{p}$. The distances $\lambda$, $L$ and $d$ are related by $\lambda\ll L\ll d$, where $\lambda$ is the characteristic wavelength of the GWs that PTAs search for (of order 10 light-years).
In transverse-traceless-synchronous coordinates $(t,\vec{x})\equiv(t,x^{i})$, where $i=1,2,3$, we have [^2] It is immediate from the functional form of $h^{A}(t-\hat{\Omega}\cdot\vec{x})$, and the definitions of the polarization tensors $e^{+}_{ij}(\hat{\Omega})$, $e^{\times}_{ij}(\hat{\Omega})$ [see (1ca), (1cb)] in terms of the unit vectors $\hat{\Omega}$, $\hat{m}$, $\hat{n}$ [see (1cda), (1cdb), (1cdc)] that the metric perturbations $h_{ij}(t,\vec{x})$ satisfy the vacuum wave equation, $(\partial^{2}/\partial t^{2}-\nabla^{2})h_{ij}=0$, are transverse $h_{ij}\hat{\Omega}^{j}=0$, and traceless $h_{ij}\delta^{ij}=0$. The coordinates are “synchronous” because the metric perturbations are purely spatial: their time components vanish.
@@
Here, $\hat{m}$ and $\hat{n}$ are any two orthogonal unit vectors in the plane orthogonal to $\hat{\Omega}$. For stochastic background analyses, a common convention which we adopt here is
@@
$\hat{n}\equiv-\hat{x}\,\cos\theta\cos\phi-\hat{y}\,\cos\theta\sin\phi+\hat{z}\,\sin\theta=-\theta\,,$ (1cdc)
@@
Figure 3: Left panel: Coordinate system and unit vectors used in the plane wave expansion. Right panel (top): two orthogonal polarizations of a GW. A circular ring of freely falling test particles is placed at rest in the plane orthogonal to the propagation direction $\hat{\Omega}$ of the wave. These are periodically deformed into ellipses, as space is “squeezed” and “stretched” by the passing of the GW. Right panel (bottom): an alternative choice of two orthogonal polarizations may be obtained by rotating $\hat{m}$ and $\hat{n}$ through angle $\psi$ to obtain new vectors $\hat{m}^{\prime}$ and $\hat{n}^{\prime}$.
@@
as can be seen from (1b). Note that this mixes $h_{+}(u)$ and $h_{\times}(u)$: $h^{\prime}_{+}$ and $h^{\prime}_{\times}$ are linear combinations of $h_{+}$ and $h_{\times}$. The gauge transformation is local, because if there are many sources with different $\hat{\Omega}$, then $\psi$ may be picked differently for each of these. The spin-two interpretation comes from quantum mechanics, where it indicates that the phase of the wave function rotates at twice the rate of the physical coordinate axes.
@@
where $A$ is a constant which is proportional to the inverse of the distance to the SMBHs. The complex waveform is
$h(t)\equiv h^{+}(t)+ih^{\times}(t)=A\exp[i(2\pi f_{0}t+\phi_{0})]\,.$ (1cdi)
Using (1b) and (1cdf), it follows that [^4] As written in (1cdj), the time dependence of the metric perturbations $h_{ij}(t,\vec{x})$ has been moved from the functions $h^{A}(t)$ into the polarization tensor $e^{\prime}{}^{+}_{ij}(\psi(t),\hat{\Omega})$, via the time dependence of $\psi(t)$. Normally, the polarization tensors are taken to be time independent.
$h_{ij}(t,\vec{0})=A\,e^{\prime}{}^{+}_{ij}(\hat{\Omega},\psi(t))\,,\quad\psi(t)\equiv\frac{1}{2}\left[2\pi f_{0}t+\phi_{0}\right]\,,$ (1cdj)
@@
Any gravitational-wave sources whose $+$ and $\times$ components satisfy (1cdk) are said to be unpolarized. At first, this seems to be a strange use of language, since it means that a circularly polarized source is “unpolarized”. A better way to think about it is this: an unpolarized source has no preferred polarization direction, as defined by the polarization angle $\psi$, in the plane perpendicular to $\hat{\Omega}$. So, according to this definition, a circularly polarized source is unpolarized, while a linearly polarized or elliptically polarized source is not unpolarized. The Hellings and Downs curve that we will discuss later is appropriate for a GWB made up of a superposition of unpolarized GW sources, or for a collection of polarized sources, provided that the collection does not define any preferred polarization angle.
@@ 2.2 Redshift response for a single Earth-pulsar baseline
Thus, the pulse “samples” the metric perturbations along its (null) path through spacetime. Note that we do not need to include any corrections to the straight-line path for the photon given above, as the metric perturbations are already first-order small and we can ignore all second- and higher-order terms in the calculation.
@@
A redshift means an apparent lowering of the clock frequency corresponding to increase in the arrival period of the pulse, while a blueshift is the opposite. While “redshift” and “blueshift” are normally used in the context of visible light, they also apply to radio signals and even to the time delays between pulses from a pulsar, since relativistic effects are felt by all types of clocks and oscillators. (A detailed derivation of this equation in the context of temperature fluctuations in the cosmic microwave background was given by Sachs and Wolfe in [13]; see Eq. (39) in that paper. This was the first such occurrence of this equation that we are aware of.)
@@
This difference between the Earth and pulsar terms has a simple physical interpretation. As explained in [14, 15], the redshift (1cdmr) can be interpreted as a Doppler shift resulting from a relative velocity between the Earth and the pulsar, projected along the line of sight to the pulsar. This projected relative velocity is obtained by parallel-transporting the pulsar 4-velocity vector along the null geodesic traversed by the radio pulse. Via the Doppler effect, this nonzero relative velocity redshifts or blueshifts the photon frequency. If both source and receiver have the same projected velocity, then the Earth and pulsar term cancel, and the GW has no net effect on the photon redshift/pulsar timing residual.
We can also write the above expression for the redshift response $Z(t)$ as
@@
At first glance it appears that the geometrical projection factor $F^{A}(\hat{\Omega})$ diverges when $\hat{\Omega}=-\hat{p}$, since the denominator $1+\hat{\Omega}\cdot\hat{p}$ in (1cdmu) then equals zero. But the reader can easily verify that this zero in the denominator is canceled by a zero from the projection $\hat{p}^{i}\hat{p}^{j}e^{A}_{ij}(\hat{\Omega})$ when $\hat{p}$ is proportional to $\hat{\Omega}$ [take $\theta\to 0$ in (1cdmak)]. In fact, the redshift response $Z(t)$ vanishes for this case because the two functions in (1cdmt) are both evaluated at $t$ when $01+\hat{\Omega}\cdot\hat{p}=$ 0. This makes sense: a GW deforms space only in the plane transverse to its direction of propagation $\hat{\Omega}$. When $\hat{p}=-\hat{\Omega}$, the radio pulse does not feel these deformations.
@@ 2.3 Transfer functions and geometric projection factors
$r^{2}=A^{2}\left|1-{\rm e}^{-i2\pi fL(1+\hat{\Omega}\cdot\hat{p})}\right|^{2}\left[F^{+}(\hat{\Omega})^{2}+F^{\times}(\hat{\Omega})^{2}\right]\,.$ (1cdmy)
@@
$r=A\bigl{|}{\cal T}(\alpha,\hat{\Omega}\cdot\hat{p})\bigr{|}|F(\Omega)|,$ (1cdmz)
@@
$|F(\Omega)|=\sqrt{F^{+}(\hat{\Omega})^{2}+F^{\times}(\hat{\Omega})^{2}}$ (1cdmaa)
is the quadrature combination of the pulsar’s responses to the two polarizations.
@@
The quadrature combination is also relevant for polarized sources, if there are no preferred orientations. To see this, first consider a single polarized source, defined by some complex waveform $h(u)=h^{+}(u)+ih^{\times}(u)$. From (1cdms), the redshift at Earth is $Z(t)=\Re[\Delta h(t,L\hat{p})F^{*}(\Omega)]$. How would that be changed if the axes which define the source’s polarization are rotated through an angle $\psi$ as shown in the bottom right panel of Figure 3? Because this does not change $\hat{\Omega}$ or the direction $\hat{p}$ to the pulsar, it only affects the redshift via $F$, which is modified because the complex polarization tensor appearing in $F$ is transformed according to (1cdf). Let $C_{R}$ and $C_{I}$ denote the real and imaginary parts of the complex quantity $C=C_{R}+iC_{I}=\Delta h(t,L\hat{p})F^{*}(\Omega)$. Then the redshift for this new source, with rotated polarization axes, is
@@
$R^{A}(f,\hat{\Omega})\simeq i2\pi fL\,G^{A}(\hat{\Omega})\,,\quad G^{A}(\hat{\Omega})\equiv\frac{1}{2}\hat{p}^{i}\hat{p}^{j}e_{ij}^{A}(\Omega)\,.$ (1cdmah)

### v3

@@ Abstract
We answer frequently asked questions (FAQs) about the Hellings and Downs correlation curve—the “smoking-gun” signature that pulsar timing arrays (PTAs) have detected gravitational waves (GWs). Many of these questions arise from inadvertently applying intuition about the effects of GWs on LIGO-like detectors to the case of pulsar timing, where not all of it applies. This is because Earth-based detectors, like LIGO and Virgo, have arms that are short (km scale) compared to the wavelengths of the GWs that they detect ($\approx 10^{2}$$10^{4}$ km). In contrast, PTAs respond to GWs whose wavelengths (tens of light-years) are much shorter than their arms (a typical PTA pulsar is hundreds to thousands of light-years from Earth). To demonstrate this, we calculate the time delay induced by a passing GW along an Earth-pulsar baseline (a “one-arm, one-way” detector) and compare it in the “short-arm” (LIGO-like) and “long-arm” (PTA) limits. This provides qualitative and quantitative answers to many questions about the Hellings and Downs curve. The resulting FAQ sheet should help in understanding the “evidence for GWs” recently announced by several PTA collaborations.
@@ 1 Introduction
These observations can be challenging to understand, in part because the observed signal in the PTA band is unlike the short duration ($\sim\!0.25~{}{\rm s}$) GW “chirp” signal that LIGO observed on 14 Sep 2015, from the final inspiral and merger of two $\sim\!30M_{\odot}$ black holes. Instead of short (1 s) audio frequency (100 Hz) deterministic GW signals, PTAs search for persistent (signals with correlation times $\lesssim 500$ yr [5, Fig. 9] and which are in band for $\sim\!10^{6}$ yr), low-frequency (nHz) stochastic GW signals. These are too feeble to observe directly, but should produce correlated perturbations in the arrival times of pulses from an array of galactic millisecond pulsars.
Although the current PTA observations are not statistically significant enough to unambiguously claim a detection or to identify a source, a credible candidate is the superposition of GWs from hundreds of thousands of supermassive ($\sim\!10^{9}M_{\odot}$) black holes (SMBH) orbiting one another in the centers of merging galaxies. While any one of those binaries would produce a deterministic almost-constant-frequency signal, the incoherent sum of the GWs from many such sources creates a stochastic confusion noise: a “hubbub” or “hum” of GWs.
PTAs search for correlations that are predicted to follow the so-called “Hellings and Downs” curve. This correlation curve is named after Ron Hellings and George Downs, who first derived it in 1983 [6]. The idea is that pulsars act as high-quality clocks, emitting radio pulses that arrive at Earth [^3] Our treatment is simplified: it assumes that the Earth is isolated, and that it is not orbiting the Sun, orbited by the Moon, or affected by other planets. In practice, these and other deterministic effects are removed by converting all pulse arrival times to pulse arrival times at the center of gravity of the solar system, which is called the Solar System Barycenter (SSB). The SSB is located near the surface of the Sun, about 8 light-minutes from Earth. On the scale of the GWs of interest, this is close enough to the Earth that we do not make a distinction here. with very predictable and regular arrival times. GWs induce slight perturbations in these arrival times, and these perturbations (called timing residuals) are correlated between different pulsars in a very specific fashion.
@@ 4. Q4:
    Why is the most negative value of the Hellings and Downs curve not at $90^{\circ}$?
@@ 8. Q8:
    Assuming noise-free observations, are the fluctuations away from the expected Hellings and Downs curve only due to the pulsar-term contributions to the timing residual response?
@@ 9. Q9:
  10. Q10:
    If a deviation from the Hellings and Downs curve is detected, what does that mean?
@@ 10. Q10:
Note that, as suggested by Q7, we make a distinction between (a) the Hellings and Downs curve, which is the black curve shown in Figure 1 and (b) the observed Hellings and Downs bin-averaged correlation, shown by the blue points and error bars in the same figure, which might or might not follow this curve.
@@ 1.2 Outline
Readers who are already familiar with pulsar timing and the definition of the Hellings and Downs correlation could skip Sections 2 and 3 and jump directly to Section 4. However, we advise against this for two reasons: (i) In Sections 2.2 and 2.4, we compare the response of “long-arm” PTA pulsars to that of “short-arm” LIGO-like detectors, which we believe is the source of many of the misconceptions regarding the Hellings and Downs curve. (ii) Our discussion of pulsar averaging in Section 3.2 will be new to many, and the concepts introduced there are important for understanding what follows.
@@
From here forward, we use units in which the propagation speed $c$ of light and of GWs is set equal to one. This means that both time and distance are measured in units of seconds or years as appropriate. More precisely, a distance is given by the time that it takes light to traverse that distance. For example, a meter stick is approximately 3 nanosec long, the LIGO detectors’ arms are about 10 microseconds long, the wavelengths of the GWs that PTAs are sensitive to are typically 10 years long, and typical pulsars are thousands of years from Earth.
@@ 2.1 Single GW point source
Figure 2: Geometry for the single pulsar analysis. The Earth is located at the origin of a rectangular Cartesian coordinate system. The components $h_{ij}(t,\vec{x})$ of the metric perturbations are functions of time and of the spatial coordinates $\vec{x}$. A single point source of GWs is located a distance $d$ from Earth in direction $-\hat{\Omega}$ on the sky. A galactic pulsar is located a distance $L$ from Earth in direction $\hat{p}$. The distances $\lambda$, $L$ and $d$ are related by $\lambda\ll L\ll d$, where $\lambda$ is the characteristic wavelength of the GWs that PTAs search for (of order 10 light-years).
In transverse-traceless-synchronous coordinates $(t,\vec{x})\equiv(t,x^{i})$, where $i=1,2,3$, we have [^4] It is immediate from the functional form of $h^{A}(t-\hat{\Omega}\cdot\vec{x})$, and the definitions of the polarization tensors $e^{+}_{ij}(\hat{\Omega})$, $e^{\times}_{ij}(\hat{\Omega})$ [see (1ca), (1cb)] in terms of the unit vectors $\hat{\Omega}$, $\hat{m}$, $\hat{n}$ [see (1cda), (1cdb), (1cdc)] that the metric perturbations $h_{ij}(t,\vec{x})$ satisfy the vacuum wave equation, $(\partial^{2}/\partial t^{2}-\nabla^{2})h_{ij}=0$, are transverse $h_{ij}\hat{\Omega}^{j}=0$, and traceless $h_{ij}\delta^{ij}=0$. The coordinates are called “synchronous” because the metric perturbations are purely spatial: their time components vanish.
@@
Here, $\hat{m}$ and $\hat{n}$ are any two orthogonal unit vectors in the plane perpendicular to $\hat{\Omega}$. For stochastic background analyses, a common convention which we adopt here is
@@
$\hat{n}\equiv-\hat{x}\,\cos\theta\cos\phi-\hat{y}\,\cos\theta\sin\phi+\hat{z}\,\sin\theta=-\hat{\theta}\,,$ (1cdc)
@@
Figure 3: Left panel: Coordinate system and unit vectors used in the plane wave expansion. Right panel (top): two orthogonal polarizations of a GW. A circular ring of freely falling test particles is placed at rest in the plane orthogonal to the propagation direction $\hat{\Omega}$ of the wave. These are periodically deformed into ellipses, as space is “squeezed” and “stretched” (here very exaggerated) by the passing of the GW. Right panel (bottom): an alternative basis of two orthogonal polarization vectors may be obtained by rotating $\hat{m}$ and $\hat{n}$ through angle $\psi$ to obtain new vectors $\hat{m}^{\prime}$ and $\hat{n}^{\prime}$.
@@
as can be seen from (1b). Note that this mixes $h^{+}(u)$ and $h^{\times}(u)$: $h^{\prime}{}^{+}$ and $h^{\prime}{}^{\times}$ are linear combinations of $h^{+}$ and $h^{\times}$. The gauge transformation is local, because if there are many sources with different $\hat{\Omega}$, then $\psi$ may be picked differently for each of these. The spin-two interpretation comes from quantum mechanics, where it indicates that the phase of the wave function rotates at twice the rate of the physical coordinate axes.
@@
where $A$ is a constant which is inversely proportional to the distance from Earth or pulsars to the SMBHs. The complex waveform is
$h(t)\equiv h^{+}(t)+ih^{\times}(t)=A\exp\left(i(2\pi f_{0}t+\phi_{0})\right)\,.$ (1cdi)
Using (1b) and (1cdf), it follows that [^6] As written in (1cdj), the time dependence of the metric perturbations $h_{ij}(t,\vec{x})$ has been moved from the functions $h^{A}(t)$ into the polarization tensor $e^{\prime}{}^{+}_{ij}(\hat{\Omega},\psi(t))$, via the time dependence of $\psi(t)$. Normally, the polarization tensors are taken to be time independent.
$h_{ij}(t,\vec{0})=A\,e^{\prime}{}^{+}_{ij}(\hat{\Omega},\psi(t))\,,\quad\psi(t)\equiv\frac{1}{2}\left(2\pi f_{0}t+\phi_{0}\right)\,,$ (1cdj)
@@
Any gravitational-wave source whose $+$ and $\times$ components satisfy (1cdk) is said to be unpolarized. At first, this seems to be a strange use of language, since it means that a circularly polarized source is “unpolarized”. A better way to think about it is this: an unpolarized source has no preferred polarization direction, as defined by the polarization angle $\psi$, in the plane perpendicular to $\hat{\Omega}$. So, according to this definition, a circularly polarized source is unpolarized, whereas a linearly polarized or elliptically polarized source is not unpolarized. The Hellings and Downs curve that we will discuss later is appropriate for a GWB made up of a superposition of unpolarized GW sources, or for a collection of polarized sources, provided that the collection does not define any preferred polarization angle.
@@ 2.2 Redshift response for a single Earth-pulsar baseline
Thus, the pulse “samples” the metric perturbations along its (null) path through spacetime. Note that we do not need to include any corrections to the straight-line path for the photon given above, as the metric perturbations are already first-order small and we can ignore all second- and higher-order terms in the calculation [13].
@@
A redshift means an apparent lowering of the clock frequency, corresponding to an increase in the arrival period of the pulse, while a blueshift is the opposite. While “redshift” and “blueshift” are normally used in the context of visible light, they also apply to radio signals and even to the time delays between pulses from a pulsar, since relativistic effects are felt by all types of clocks and oscillators. (A detailed derivation of this equation in the context of temperature fluctuations in the cosmic microwave background was given by Sachs and Wolfe in [13]; see Eq. (39) in that paper. This was the first such occurrence of this equation that we are aware of.)
@@
This difference between the Earth and pulsar terms has a simple physical interpretation. As explained in [14, 15, 16], the redshift (1cdmr) can be interpreted as a Doppler shift resulting from a relative velocity between the Earth and the pulsar, projected along the line of sight to the pulsar. This projected relative velocity is obtained by parallel-transporting the pulsar 4-velocity vector along the null geodesic traversed by the radio pulse. Via the Doppler effect, this nonzero relative velocity redshifts or blueshifts the photon frequency. If both source and receiver have the same projected velocity, then the Earth and pulsar term cancel, and the GW has no net effect on the photon redshift/pulsar timing residual.
We can write the above expression for the redshift response $Z(t)$ as
@@
At first glance it appears that the geometrical projection factor $F^{A}(\hat{\Omega})$ diverges if the GW source is located behind the pulsar ($\hat{\Omega}=-\hat{p}$), since the denominator $1+\hat{\Omega}\cdot\hat{p}$ in (1cdmu) then vanishes. But the reader can easily verify that this zero in the denominator is canceled by a zero from the projection $\hat{p}^{i}\hat{p}^{j}e^{A}_{ij}(\hat{\Omega})$ when $\hat{p}$ is proportional to $\hat{\Omega}$. In this $\hat{\Omega}\to-\hat{p}$ case, the response $|F|\to 1$ is finite, as can be seen clearly from (1cdmaa). Nevertheless, the redshift $Z(t)$ vanishes: because $01+\hat{\Omega}\cdot\hat{p}=$ 0, the two functions in (1cdmt) are both evaluated at $t$, and cancel. Physically, the radio pulse “surfs” on the GW, so the Earth and pulsar term are identical and their difference is zero. However, if the GW direction $\hat{\Omega}$ is displaced slightly away from $-\hat{p}$, the response has its maximum value as shown by the nearly vertical “lobes” in the rightmost panel of Figure 6.
@@ 2.3 Transfer functions and geometric projection factors
$r^{2}=A^{2}\left|1-{\rm e}^{-i2\pi fL(1+\hat{\Omega}\cdot\hat{p})}\right|^{2}\ |F(\hat{\Omega})|^{2}$ (1cdmy)
@@
$r=A\bigl{|}{\cal T}(\alpha,\hat{\Omega}\cdot\hat{p})\bigr{|}|F(\hat{\Omega})|,$ (1cdmz)
@@
$|F(\hat{\Omega})|\equiv\sqrt{F^{+}(\hat{\Omega})^{2}+F^{\times}(\hat{\Omega})^{2}}=\frac{1}{2}(1-\hat{\Omega}\cdot\hat{p})$ (1cdmaa)
is the quadrature combination of the pulsar’s responses to the two polarizations. We recommend that the reader verify the last equality in (1cdmaa); see also [17, Eq. (2.7)].
@@
The quadrature combination is also relevant for polarized sources, if there are no preferred orientations. To see this, first consider a single polarized source, defined by some complex waveform $h(u)=h^{+}(u)+ih^{\times}(u)$. From (1cdms), the redshift at Earth is $Z(t)=\Re[\Delta h(t,L\hat{p})F^{*}(\hat{\Omega})]$. How would that be changed if the axes which define the source’s polarization are rotated through an angle $\psi$ as shown in the bottom right panel of Figure 3? Because this does not change $\hat{\Omega}$ or the direction $\hat{p}$ to the pulsar, it only affects the redshift via $F$, which is modified because the complex polarization tensor appearing in $F$ is transformed according to (1cdf). Let $C_{R}$ and $C_{I}$ denote the real and imaginary parts of the complex quantity $C=C_{R}+iC_{I}=\Delta h(t,L\hat{p})F^{*}(\hat{\Omega})$. Then the redshift for this new source, with rotated polarization axes, is
@@
$R^{A}(f,\hat{\Omega})\simeq i2\pi fL\,G^{A}(\hat{\Omega})\,,\quad G^{A}(\hat{\Omega})\equiv\frac{1}{2}\hat{p}^{i}\hat{p}^{j}e_{ij}^{A}(\hat{\Omega})\,.$ (1cdmah)